# Exploring Decision-based Black-box Attacks on Face Forgery Detection

### v1

@@ IV-B Attacks on State-of-the-art Detectors
Attack on Spatial-based Models. Table II shows the attack success rates against spatial-based models on FaceForensics++ and CelebDF. Compared with other attacks, our method has a higher query efficiency on the premise of ensuring 100% attack success rate. Furthermore, Boundary, SignFlip, and RayS fail to attack some models (e.g. ResNet50, Eb4) even with real faces as perturbation initialization. For example, SignFlip and RayS attack Eb4 fails on FaceForensics++, but our method achieves the attack with a high attack efficiency. Meanwhile, compared to SignFlip and RayS, the average query number of our method is 1/2 and 1/7 of that on ResNet50, and 1/5 and 1/160 of that on Xception. On CelebDF, our method improves the attack efficiency by 17.90% and 23.81% against Xception and Eb4 compared to the state-of-the-art SignFlip. Although less related to the features of ViT, our method can also achieve a state-of-the-art attack success rate, which effectively illustrates the generalization of our method.
@@
TABLE IV: Ablation study on key modules. Frequency noises and cross-task perturbation initialization significantly improve the attack efficiency, illustrating its effectiveness.
@@
[…]
TABLE V: The selection of intermediate feature layers of face recognition models. The bolded intermediate feature layers are used to implement cross-task perturbation initialization.
@@
IV-C Ablation Study
In this section, we choose ResNet50 [1] and ViT [4] as the threat models and randomly select 128 faces (4$\times$32=128) on FaceForensics++ [18] for the ablation study.
Key Modules. We first investigate the impact of the proposed key modules on the attack efficiency. Fre_Noise represents whether to add noises in the frequency domain (refer to Eq. 3), and CPI represents cross-task perturbation initialization. As shown in Table IV, adding frequency noises or cross-task perturbation initialization individually brings gains (i.e., 6.65%/52.84% on ResNet50, 2.23%/4.73% on ViT), revealing the effectiveness of the corresponding key modules. Combing all modules together leads to the best attack performance, yielding the query efficiency of 58.90% and 5.69% on ResNet50 and ViT, respectively.
Intermediate Feature Layers. We choose different face recognition models (FaceNet [47], CosFace[49], and ArcFace[50]) and consider choosing different intermediate feature layers to craft cross-task adversarial perturbations. Table V illustrates the selection of intermediate feature layers, and Fig. 4 illustrates the average number of queries for different intermediate feature layers. Attacking the features of the middle layer achieves better results than attacking the top or bottom layers, so we choose to use the bold layers in Table V to generate cross-task adversarial perturbations. Besides, FaceNet has better attack efficiency than other models, so we choose FaceNet as the face recognition model initialized by cross-task perturbation.
TABLE VI: Ablation study on the frequency noise magnitude $\gamma$.
@@ IV-C Ablation Study
Frequency noise magnitude $\gamma$. The magnitude of the frequency perturbation is critical to the efficiency of the attack. Table VI illustrates the effect of different frequency noise magnitudes $\gamma$ on the attack efficiency. When the $\gamma$ is $1.75$, ResNet50 and ViT achieve the smallest AQ and MQ, and maintain a 100% attack success rate. Therefore, the $\gamma$ in this paper is $1.75$.
@@
IV-D Image Quality Assessment
In this section, we analyze the image quality assessment of adversarial examples generated on FaceForensics++ [18]. Due to page limitations, here we choose Xception as the threat model, and state-of-the-art Signflip [33] and RayS [35] for comparison. To assess the image quality, MSE, PSNR, and SSIM [51] are used as evaluation metrics to quantify the difference between adversarial and real faces. As shown in Table VII, our method outperforms other attacks by a wide margin in terms of MSE, PSNR, and SSIM. Further, we visualize the adversarial examples generated by SignFlip, RayS, and our method in Fig. 5. The adversarial examples generated by our method are more imperceptible, unlike SignFlip with artifacts or RayS with obvious black lines. Finally, to verify that better image quality can attack both face recognition and face forgery detection, we choose the recognition success rate (RSR, %) as the evaluation criterion. We compare the face of adversarial examples in Table II with the real face of its same identity, and the results are shown in Table VII. In addition, frequency noise significantly improves image quality. It shows that our method has stronger attack performance while maintaining high image quality since our method is able to break both face recognition and face forgery detection.
IV-E Attacks on industrial APIs
In this section, we investigate the applicability of decision-based attacks on real-world systems, such as commercial face forgery detection APIs. We regard the face forgery detection API [^2] https://cloud.tencent.com/product/atdf?fromSource=gwzcw on Tencent AI Open Platform as the threat model. We choose the detection threshold to be 0.5, that is, if the output is greater than the threshold, the input is judged as fake faces. Simultaneously, we choose the face recognition API [^3] https://ai.qq.com/product/face.shtml#compare on Tencent AI Open Platform to calculate the recognition success rate (RSR, %). For face comparison, we choose a similarity threshold of 60. We choose 100 random frames of FaceForensics++ [18] as test samples, including four forging methods, Face2Face [10], FaceSwap [11], DeepFakes [12] and NeuralTextures [13].
Here, we set the maximum number of queries to $1,000$ and the maximum perturbation $\epsilon$ to 0.062 and 0.125. RSR is only calculated on successful samples identified as real faces. Table VIII shows the experimental results of attacking industrial APIs, and Figure 6 shows the visualization of attacking commercial APIs. Note that RayS [35] is ineffective in attacking industrial APIs. This is because RayS’ perturbation initialization causes the image to become black, making the API unable to detect faces, and eventually leading to API denial of service. Compared with SignFlip [33], our method has a higher attack success rate (26% $\uparrow$ and 1% $\uparrow$) and lower query efficiency under different $\epsilon$. In terms of recognition success rate (RSR), our method outperforms the baseline with an improvement of 17.00% and 3.00%, respectively, which can be attributed to the better image quality produced by our method. This suggests that our method is more effective in simultaneously attacking face recognition and face forgery detection systems. Therefore, our method can effectively evaluate the security of existing AI systems in real-world scenarios.
TABLE VII: Quantitative image quality assessment of adversarial examples generated by SignFlip [33], RayS [35], and our method on FaceForensics++ [18].
@@ IV-E Attacks on industrial APIs
TABLE VIII: Experimental results for attacking industrial APIs.
@@
IV-F Adverarial Transferability
In this section, we evaluate the cross-dataset adversarial transferability of different decision-based attacks. We generate adversarial examples on CelebDF [19] and then test the same model on FFDF [18]. Table IX shows that the attack success rate (%) on cross-dataset evaluation. Because our method can eliminate the difference between real and fake faces in the frequency domain, it has stronger cross-dataset adversarial transferability.
TABLE IX: Cross-dataset evaluation on adversarial transferability.
@@ IV-F Adverarial Transferability
IV-G JPEG Defense
JPEG [52] is a lossy image compression method designed to preferentially preserve details important to the human visual system, and therefore has become a mainstream adversarial defense method. To further evaluate the robustness of our method, we evaluate JPEG defenses in the form of adaptive attacks [52]. As shown in Table X, since JPEG [52] removes high-frequency features, our FDA adds noise at both high and low frequencies, thus marginally reducing attack efficiency but keeping 100% ASR because high-frequency noise is still effective.
TABLE X: Attack performance with the JPEG defense on FFDF.
[…]
Ours 21.70 12.0 100.00
Ours+JPEG 36.88 26.0 100.00
V Discussion
In this section, we further analyze the activation relationship of Cross-task Perturbation Initialization in face recognition and face forgery detection features in Section III-B. Face recognition distinguishes the identities of different faces. Recent work [53, 54] focuses on using face identities to enhance face forgery detection. The middle layer features of face recognition are positively correlated with face forgery detection, so the adversarial examples of face recognition can be used to attack face forgery detection, and we propose Cross-task Perturbation Initialization. Our work is the first to simultaneously attack two face-related tasks, face forgery detection and face recognition, in academic and industrial scenarios. Cross-task Perturbation Initialization on face recognition is also beneficial to attack face recognition and face forgery detection. Therefore, the introduction of face recognition may be the future research direction of face forgery detection.
VI Conclusions
In this paper, we first explore decision-based attacks on face forgery detection. The direct application of existing decision-based attacks, however, suffers from initialization failures and low image quality. To alleviate initialization sensitivity on attacks, we propose cross-task perturbation that utilizes the high correlation of face features across tasks. Based on the frequency cues utilized in face forgery detection, we propose a frequency decision-based attack. In particular, we perturb the frequency domain and then constrain the visual quality in the spatial domain. Our method achieves high query efficiency and guaranteed image quality for FaceForensics++, CelebDF, and APIs, demonstrating state-of-the-art attack performance.

### v2

@@ 4.2. Attacks on State-of-the-art Detectors
Attack on Spatial-based Models. Table 2 shows the attack success rates against spatial-based models on FaceForensics++ and CelebDF. Compared with other attacks, our method has a higher query efficiency on the premise of ensuring 100% attack success rate. Furthermore, Boundary, SignFlip, and RayS fail to attack some models (e.g. ResNet50, Eb4) even with real faces as perturbation initialization. For example, SignFlip and RayS attack Eb4 fails on FaceForensics++, but our method achieves the attack with a high attack efficiency. Meanwhile, compared to SignFlip and RayS, the average query number of our method is about 1/2 and 1/4 of that on ResNet50, and 1/5 and 1/160 of that on Xception. On CelebDF, our method improves the attack efficiency by 17.90% and 23.81% against Xception and Eb4 compared to the state-of-the-art SignFlip. Although less related to the features of ViT, our method can also achieve a state-of-the-art attack success rate, which effectively illustrates the generalization of our method.
Attack on Frequency-based Models. Face forgery detection often uses the difference in frequency as a clue to judge whether it is forgery. To verify the adversarial robustness of these frequency-based methods, we choose F3Net [f3net] and M2TR [M2TR] for experiments. Table 3 shows the attack success rates of frequency-based detectors on FaceForensics++ and CelebDF. Frequency-based detectors perform well on face forgery detection, but they are just as susceptible to adversarial robustness as spatial-based detectors. On F3Net, our method achieves state-of-the-art attack performance, especially on CelebDF, with 35.56% fewer queries on average than SignFlip. On M2TR, our method improves the query efficiency by 22.41% and 61.20% over state-of-the-art methods on FaceForensics++ and CelebDF, respectively. Although frequency-based face forgery detectors have better detection performance than spatial-based detectors, they are both vulnerable to adversarial examples. Our attack adds noise in the frequency domain and has high efficiency in attacking both spatial and frequency domain based detectors, which illustrates the effectiveness of frequency noise. Therefore, the above experiments demonstrate the effectiveness of our method and further expose the adversarial vulnerability of face forgery detection.
Attack on Self-blended Models. [sbi] presents an innovative deepfake detection method utilizing self-blended images (SBIs), which are artificially generated by blending pseudo source and target images derived from a single authentic image. This approach replicates typical forgery artifacts, such as blending boundaries and statistical discrepancies, thus producing more generalized and less distinguishable fake samples. Table 5 shows the attack performance of SBI [^4] We use the implementation and weights (FF-c23) provided by https://github.com/mapooon/SelfBlendedImages. on FaceForensics++ [ffpp]. Except for Rays and Ours, other methods use target images as initialization. For SBI, these are obviously blended images, so they will always be identified as fake faces, resulting in unsuccessful attacks. Due to the efficient initialization provided by CPI, our attack maintains a 100% attack success rate (ASR) on SBI, while requiring only approximately 28.33% of the average queries needed by RayS. This demonstrates the continued effectiveness of our method, even against more generalized face forgery detectors.
Attack on different perturbations $\epsilon$. To further improve our adversarial robustness evaluation, we test the attack performance with different maximum perturbations $\epsilon$ on FaceForensics++’s ResNet and ViT-B, as shown in Table 4. We evaluate the attack at $\epsilon$ of 0.031, 0.05, and 0.062, respectively, and find that the less the maximum perturbation, the lower the attack efficiency. Especially for HSJA, on ViT-B, when $\epsilon=0.062$, the AQ and ASR are 264.23 and 91.61. When $\epsilon=0.031$, the AQ increases slightly to 288.08, but the ASR drops significantly to 38.03. SignFlip, RayS, and Ours have high ASR at different eps, but our attack has the best attack efficiency, which shows its applicability in adversarial robustness evaluation.
4.3. Ablation Study
[…]
@@ 4.3. Ablation Study
Table 6: Ablation study on key modules. Frequency noises and cross-task perturbation initialization significantly improve the attack efficiency, illustrating its effectiveness.
@@
Key Modules. We first investigate the impact of the proposed key modules on the attack efficiency. Fre_Noise represents whether to add noises in the frequency domain (refer to Eq. 3), and CPI represents cross-task perturbation initialization. As shown in Table 6, adding frequency noises or cross-task perturbation initialization individually brings gains (i.e., 6.65%/52.84% on ResNet50, 2.23%/4.73% on ViT), revealing the effectiveness of the corresponding key modules. Combing all modules together leads to the best attack performance, yielding the query efficiency of 58.90% and 5.69% on ResNet50 and ViT, respectively.
Intermediate Feature Layers. We choose different face recognition models (FaceNet [FaceNet] [^5] https://github.com/ShawnXYang/Face-Robustness-Benchmark/blob/master/RobFR/networks/FaceNet.py, CosFace[CosFace] [^6] https://github.com/ShawnXYang/Face-Robustness-Benchmark/blob/master/RobFR/networks/CosFace.py, and ArcFace[ArcFace] [^7] https://github.com/ShawnXYang/Face-Robustness-Benchmark/blob/master/RobFR/networks/ArcFace.py) and consider choosing different intermediate feature layers to craft cross-task adversarial perturbations. Table 7 illustrates the selection of intermediate feature layers, and Figure 4 illustrates the average number of queries for different intermediate feature layers. Attacking the features of the middle layer achieves better results than attacking the top or bottom layers, so we choose to use the bold layers in Table 7 to generate cross-task adversarial perturbations. Besides, FaceNet has better attack efficiency than other models, so we choose FaceNet as the face recognition model initialized by cross-task perturbation.
Table 7: The selection of intermediate feature layers of face recognition models. The bolded intermediate feature layers are used to implement cross-task perturbation initialization.
@@
Table 8: Ablation study on the frequency noise magnitude $\gamma$.
@@
Frequency noise magnitude $\gamma$. The magnitude of the frequency perturbation is critical to the efficiency of the attack. Table 8 illustrates the effect of different frequency noise magnitudes $\gamma$ on the attack efficiency. The hyperparameter $\gamma$ is set to 1.75 based on a comprehensive analysis of ablation study results. This value achieves an optimal balance among multiple key metrics across ResNet50 and ViT. Specifically, compared to larger $\gamma$ values (e.g., 2.0 and 2.5), a $\gamma$ of 1.75 reduces AQ for ViT from 171.29 to 158.42, indicating improved efficiency, while maintaining comparable AQ for ResNet50. Meanwhile, decreasing $\gamma$ below 1.75 (e.g., 1.5 and 1.0) increases MQ for ViT. Thus, $\gamma=1.75$ optimizes both efficiency and stability, and demonstrates generalizability across different model architectures, making it the ideal choice for practical deployment.
@@
4.4. Image Quality Assessment
In this section, we analyze the image quality assessment of adversarial examples generated on FaceForensics++ [ffpp]. We choose Xception as the threat model, and state-of-the-art Signflip [SignFlip] and RayS [RayS] for comparison. To assess the image quality, MSE, PSNR, and SSIM [ssim] are used as evaluation metrics to quantify the difference between adversarial and real faces. As shown in Table 9, our method outperforms other attacks by a wide margin in terms of MSE, PSNR, and SSIM. Further, we visualize the adversarial examples generated by SignFlip, RayS, and our method in Figure 5. The adversarial examples generated by our method are more imperceptible, unlike SignFlip with artifacts or RayS with obvious black lines. Finally, to verify that better image quality can attack both face recognition and face forgery detection, we choose the recognition success rate (RSR, %) as the evaluation criterion. We compare the face of adversarial examples in Table 2 with the real face of its same identity, and the results are shown in Table 9. In addition, frequency noise significantly improves image quality. It shows that our method has stronger attack performance while maintaining high image quality since our method is able to break both face recognition and face forgery detection.
Table 9: Quantitative image quality assessment of adversarial examples generated by SignFlip [SignFlip], RayS [RayS], and our method on FaceForensics++ [ffpp].
@@ 4.4. Image Quality Assessment
4.5. Attacks on industrial APIs
In this section, we investigate the applicability of decision-based attacks on real-world systems, such as commercial face forgery detection APIs. We regard the face forgery detection API [^8] https://cloud.tencent.com/product/atdf?fromSource=gwzcw on Tencent AI Open Platform as the threat model. We choose the detection threshold to be 0.5, that is, if the output is greater than the threshold, the input is judged as fake faces. Simultaneously, we choose the face recognition API [^9] https://ai.qq.com/product/face.shtml##compare on Tencent AI Open Platform to calculate the recognition success rate (RSR, %). For face comparison, we choose a similarity threshold of 60. We choose 100 random frames of FaceForensics++ [ffpp] as test samples, including four forging methods, Face2Face [Face2Face], FaceSwap [Faceswap], DeepFakes [Deepfake] and NeuralTextures [neuraltextures].
Here, we set the maximum number of queries to $1,000$ and the maximum perturbation $\epsilon$ to 0.062 and 0.125. RSR is only calculated on successful samples identified as real faces. Table 10 shows the experimental results of attacking industrial APIs, and Figure 6 shows the visualization of attacking commercial APIs. Note that RayS [RayS] is ineffective in attacking industrial APIs. This is because RayS’ perturbation initialization causes the image to become black, making the API unable to detect faces, and eventually leading to API denial of service. Compared with SignFlip [SignFlip], our method has a higher attack success rate (26% $\uparrow$ and 1% $\uparrow$) and lower query efficiency under different $\epsilon$. In terms of recognition success rate (RSR), our method outperforms the baseline with an improvement of 17.00% and 3.00%, respectively, which can be attributed to the better image quality produced by our method. This suggests that our method is more effective in simultaneously attacking face recognition and face forgery detection systems. Therefore, our method can effectively evaluate the security of existing AI systems in real-world scenarios.
Table 10: Experimental results for attacking industrial APIs.
@@ 4.5. Attacks on industrial APIs
4.6. Adversarial Transferability
Our approach is query-based, achieving an approximate 100% attack success rate with exceptionally high query efficiency. To further enhance our evaluation, we assess the cross-dataset adversarial transferability of various decision-based attacks. Specifically, we generate adversarial examples on the CelebDF dataset [CelebDF] and subsequently test the same model on the FFDF dataset [ffpp]. Table 11 illustrates the attack success rate (%) in this cross-dataset evaluation. Our method demonstrates superior cross-dataset adversarial transferability due to its ability to mitigate differences between real and fake faces in the frequency domain.
Table 11: Cross-dataset evaluation on adversarial transferability.
@@ 4.6. Adversarial Transferability
4.7. JPEG Defense
JPEG [jpeg] is a lossy image compression method designed to preferentially preserve details important to the human visual system. The human eye is more sensitive to low-frequency components and less sensitive to high-frequency details. JPEG compression takes advantage of this visual characteristic and can significantly reduce the file size while maintaining the visual quality of the image by removing or reducing the amount of images of high-frequency components. Therefore, it has become a mainstream adversarial defense method to eliminate noises. To further evaluate the robustness of our attack, we evaluate JPEG defenses [jpeg] under different compression rates. As shown in Table 12, a stronger compression rate reduces the detection accuracy of the image. Since JPEG [jpeg] removes high-frequency features, our FDA adds noise at both high and low frequencies, thus keeping state-of-the-art attack performance because low-frequency noise is still effective. In addition, the attack performance of SignFlip and RayS is degraded because they do not consider the frequency noise, which shows their shortcomings in robustness evaluation.
Table 12: Attack performance with the JPEG defense on FFDF.
Attack w.o. defense JPEG Compression Rate
0.5 0.7 0.9
Clean 0.00 24.82 6.61 0.17
SignFlip 100.00 92.14 83.21 86.07
RayS 100.00 95.16 93.38 81.75
Ours 100.00 95.16 95.54 95.00
5. Discussion
In this section, we further analyze the activation relationship of Cross-task Perturbation Initialization in face recognition and face forgery detection features in Section 3.2. Face recognition distinguishes the identities of different faces. Recent work [zhuang2022towards, yu2023diff] focuses on using face identities to enhance face forgery detection. The middle layer features of face recognition are positively correlated with face forgery detection, so the adversarial examples of face recognition can be used to attack face forgery detection, and we propose Cross-task Perturbation Initialization. Our work is the first to simultaneously attack two face-related tasks, face forgery detection and face recognition, in academic and industrial scenarios. Cross-task Perturbation Initialization on face recognition is also beneficial to attack face recognition and face forgery detection. Therefore, the introduction of face recognition may be the future research direction of face forgery detection.
@@ 5. Discussion
6. Conclusions
In this paper, we first explore decision-based attacks on face forgery detection. The direct application of existing decision-based attacks, however, suffers from initialization failures and low image quality. To alleviate initialization sensitivity on attacks, we propose cross-task perturbation that utilizes the high correlation of face features across tasks. Based on the frequency cues utilized in face forgery detection, we propose a frequency decision-based attack. In particular, we perturb the frequency domain and then constrain the visual quality in the spatial domain. Our method achieves high query efficiency and guaranteed image quality for FaceForensics++, CelebDF, and industrial APIs, demonstrating state-of-the-art attack performance and exposing the adversarial vulnerability of existing detectors. We hope that this method can better evaluate the adversarial robustness of face forgery detection.
CRediT authorship contribution statement
Zhaoyu Chen: Conceptualization, Investigation, Experimentation, Visualization, Writing-Original Draft, Review & Editing. Bo Li: Supervision, Writing-Original Draft, Review & Editing. Kaixun Jiang: Experimentation, Review & Editing. Shuang Wu: Investigation, Review & Editing. Shouhong Ding: Review & Editing Wenqiang Zhang: Supervision, Review & Editing, Funding acquisition
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No.62072112), Scientific and Technological innovation action plan of Shanghai Science and Technology Committee (No.22511102202), Fudan Double First-class Construction Fund (No. XM03211178). This work was completed by Zhaoyu Chen during his internship at Youtu Lab, Tencent.